# LLM-hRIC: LLM-empowered Hierarchical RAN Intelligent Control for O-RAN

### v2

@@ Abstract
Index Terms:
Open radio access network, large language model, RAN intelligent control, reinforcement learning
@@ I Introduction
Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8]. Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4]. The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location. Furthermore, the cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides the high-level guidance to near-RT RIC, including the network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then refines and implements these policies based on real-time information to enhance the network performance.
Meanwhile, the LLMs have opened new routes for wireless network design and operation including O-RAN. The LLMs are advanced ML models, designed to understand and generate human language. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Recent studies have applied LLM-based technologies to wireless network for the resource allocation[9, 10], base station placement[11, 12], and packet analysis [13]. Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
Despite these advancements in RICs and LLM-based techniques for wireless networks, critical challenges still remain as follows: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
@@ II LLM-empowerd hierarchical RIC Framework
In this section, we introduce the LLM-empowered hierarchical RIC (LLM-hRIC) framework for resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view (i.e., local information of RAN) to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Structure of the proposed LLM-hRIC framework.
@@ II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multimodal data including text, images, and signals, enabling to form of a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
The process of the LLM-empowered non-RT RIC can be summarized as follows.
@@
    Data Integration: The non-RT RIC collects multimodal data (e.g., textual report on network conditions and radio spectrum images) and integrates the collected data by structuring it into a unified format.
@@
    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate strategic guidance (e.g., transmit power allocation strategies), which can be helpful to make decisions in the near-RT RIC (e.g., RL-based power control). The mathematical formulation of the guidance generation process is given as
    $\textbf{y}=\text{LLM}_{\theta}(x_{1},x_{2},\cdots,x_{N}),$ (1)
    where $x_{n}\in\mathcal{X}$ represents a specific data modality in the input set $\mathcal{X}$ and $\theta$ denotes the parameters of LLM.
@@ II-B RL-Empowered near-RT RIC
The near-RT RIC operates as a political layer of O-RAN, managing and optimizing network operations in near-RT. To achieve this, RL-based xApps can be deployed for near-RT RIC to determine optimal actions based on 1) the observed local information $\boldsymbol{o}$ via E2 interface and 2) the guidance vector y, received from the LLM-empowered non-RT RIC via A1 interface. Specifically, the optimal action $\boldsymbol{a}$ can be determined as
$\boldsymbol{a}=\text{RL}_{\phi}(\textbf{y},\boldsymbol{o}),$ (2)
where $\phi$ represents the parameters of the RL model.
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the RL-based xApps need to be designed to process information efficiently and make a decision fast. From the integration of LLM-empowered non-RT RIC, the RL component’s ability to refine its decision-making can be enhanced due to the LLM-generated insights (i.e., guidances), obtained from the information on large-scale network environments. This can enable the RL-based xApps to decide with a large-scale view of the network (e.g., overall traffic distribution and inter-cell interference) as well as the localized and real-time observation (e.g., instantaneous user demand and channel quality) of the network.
@@ III Use Case: Power Allocation in IAB Network
In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework.
@@ III-A Power Allocation in IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as =$\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users $\mathcal{K}_{m,n}$, connected with the $n$-th SBS of the $m$-th MBS, follow a Gaussian-Markov mobility model. The SBS allocates fixed power to their connected user. The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
@@ III-C LLM-assisted Three Phases Training for RL
  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$, to encourage exploration in the vicinity of the LLM’s guidance.
  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 00 over time.
@@
In each training stage, the power allocation policy interacts with environment, and the resulting MDP experiences are collected to train the RL agent.
@@ III-D Experimental Results
In this simulation, Llama-3.1-8B-instruct model [15] is deployed in the rApp to generate a power allocation guidance policy in a large time scale (e.g., 2s). Meanwhile, a deep deterministic policy gradient (DDPG) agent is deployed in the xApp to provide fine-grained power allocation policy for a MBS in a small time scale (e.g., 200ms). To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 3: Training curve for proposed and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
Figure 3 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$. The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies. Notably, the proposed LLM-hRIC framework achieves higher total throughput compared to the baseline methods. This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance.
@@
Figure 4 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput. The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases. This is because the performance bottleneck exists at the backhaul link for a low $\alpha$, while it is primarily in the access link for large $\alpha$. The proposed LLM-hRIC framework consistently outperforms baseline methods. This is due to the high-quality initial policies, generated by LLM, which enhances exploration efficiency and facilitates better policy discovery by RL. The LLM leverage detailed, and global information to improve coordination among MBSs, resulting in a more efficient resource allocation and higher overall network performance.
@@ IV-B Fine-Tuning of LLMs
Current LLMs are trained for general domains but not optimized for wireless communication tasks, limiting guidance quality in LLM-hRIC. While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size. Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC. However, two challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging. 2) Generating high-quality instruction data for fine-tuning is essential. Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs.
@@ V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores guidance with local observation to facilitate near-RT decision-making. The superior performance highlight the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems. Additionally, we identify key challenges for LLM-hRIC, as discussed in Section IV, to guide future research and development.

### v3

# arXiv:2504.18062v3[cs.NI] 27 Feb 2026

@@ I Introduction
Recently, intensive researches on RICs have been appeared[3, 4, 5, 6, 7, 8, 9]. Specifically, the near-RT RICs has been developed to manage a smaller number of devices and optimize on a small-time scale, such as the radio card switching for energy efficiency [3] and the flow-split distribution, congestion control, and scheduling for higher network utility [4]. In particular, deep reinforcement learning (DRL) has emerged as a key technology for intelligent xApps, enabling dynamic control of radio resource management in O-RAN [9]. The non-RT RIC has been developed to manage a larger number of devices and perform high-level and long-term network management, such as the orchestRAN[5], which automatically computes the optimal set of algorithms and their execution location. Furthermore, the cooperation between non-RT RIC and near-RT RIC is investigated in [6, 7, 8]. Specifically, the non-RT RIC provides the high-level guidance to near-RT RIC, including the network configuration and reconfiguration information [6], control signals [7], and optimized control functions [8]. The near-RT RIC then refines and implements these policies based on real-time information to enhance the network performance.
Meanwhile, the LLMs have opened new routes for wireless network design and operation including O-RAN. The LLMs are advanced ML models, designed to understand and generate human language. Their ability to understand context and generate contextually relevant outputs makes them highly adaptable across various domains. In the context of telecommunications, LLMs offer unique advantages by processing and analyzing extensive network data to improve decision-making. Recent studies have applied LLM-based technologies to wireless network for the resource allocation[10, 11], base station placement[12], and packet analysis [13]. Recently, the LLM-based xApp to allocate physical resource block (PRB) for O-RAN has also been introduced[14].
Despite advancements in RICs and LLM-based techniques for wireless networks, a critical real-world problem remains: bridging the ‘coordination gap’ between the strategic, long-term control of the non-RT RIC and the tactical, real-time execution of the near-RT RIC. This gap is exacerbated by several underlying challenges: 1) traditional ML techniques struggle to process the diverse and complex information in O-RAN; 2) the efficient cooperation between non-RT RIC and near-RT RIC still remains largely unexplored; 3) although the potential on the role of LLMs in wireless networks has been shown, the high computational complexity hinders near-RT decision-making; and 4) without fine-tuning for domain-specific tasks, the performance of LLM-based techniques often remains comparable to or worse than traditional algorithms.
@@ II LLM-empowerd hierarchical RIC Framework
In this section, we introduce the LLM-empowered hierarchical RIC (LLM-hRIC) framework for resource management in O-RAN as shown in Fig. 1. The LLM-hRIC framework consists of 1) the LLM-empowered non-RT RIC (i.e., upper layer control) and 2) the RL-empowered near-RT RIC (i.e., lower layer control), and this shows how the LLM can be used as a way to facilitate the resource management in O-RAN. As the LLM has the advantages of generalization across various environments with minimal language data, it is used for the non-RT RIC to provide guidance on the strategies to the near-RT RIC from the large-scale view of the networks. The near-RT RIC can be designed to utilize this guidance in conjunction with the small-scale view to make decisions in near-RT. More details on the LLM-hRIC framework are provided in the following subsections.
Figure 1: Structure of the proposed LLM-hRIC framework and time diagrame.
@@ II-A LLM-Empowered non-RT RIC
The non-RT RIC serves as a strategic layer, which works based on the collected KPMs from various network units including gNB and near-RT RIC via O1 interface. The rApps, deployed for the non-RT RIC, are responsible for utilizing that information to generate strategic guidance for the near-RT RIC. By integrating LLM technology, the non-RT RIC gains the capability to process and analyze multi-modal data to form a comprehensive understanding of the current network environments. Additionally, LLMs excels in synthesizing diverse data sources and leveraging their extensive parameters and deep architecture to perform complex reasoning and inference. These strengths allow the LLM-empowered non-RT RIC to tackle challenges in complex network management and provide high-quality strategic guidance to the near-RT RIC.
The process of the LLM-empowered non-RT RIC can be summarized as following steps.
@@
    Data Integration: The non-RT RIC ingests multi-modal data from gNB and near-RT RIC, primarily via the O1 interface. These data include long-term time-series metrics (e.g., total throughput trends), unstructured textual data (e.g., fault alarm logs), topology data representing network node connectivity, and even advanced visual data such as radio maps generated from digital twins.
@@
    Input Validation and Prompt Construction: Before being formatted into a prompt, this data undergoes a validation stage to handle potentially noisy or corrupted measurements. The data integration process involves structuring these varied inputs into a unified, context-rich format that can be effectively processed by the LLM.
@@
    Environment Understanding and Actionable Insights Generation: The LLM analyzes and understands the context based on the integrated multimodal data, and generate the text response, that includes the strategic guidance y (e.g., transmit power allocation strategies).
  • 4)
    Guidance Extraction and Verification: The LLM’s raw text output is validated for its conformance to predefined formats and constraints, and extracts the strategic guidance y. To ensure robustness, if this process fails, a fallback mechanism provides a default policy as guidance.
  • 5)
@@ II-B RL-Empowered near-RT RIC
Generally, given the strict latency constraints of the near-RT RIC (e.g., 10 ms to 1s), the xApps need to be designed to process information efficiently and make fast decisions in a highly dynamic network environment. Hence, we employ RL because its ability to learn an adaptive policy is well-suited for reacting to the dynamic environments with real-time changes. Furthermore, RL is naturally designed to incorporate a strategic guidance vector as a part of its state input, fusing it with the observed local information to make refined and context-aware decisions.
The process of the RL-empowered near-RT RIC can be summarized in the following steps:
  • 6)
    Data Integration: The near-RT RIC ingests the strategic guidance from the non-RT RIC (via the A1 interface) and the observed local information from the gNB (via the E2 interface).
  • 7)
    Policy generation: The RL-based xApp then analyzes these inputs to generate an optimal, fine-grained policy that is sent to the gNB for execution.
This integration of global insights with real-time local information enhances the RL agent’s ability to make refined, fine-grained decisions that are both strategically sound and tactically optimal for the immediate radio conditions.
@@ III Use Case: Power Allocation in IAB Network
In this section, we present a use case based on IAB networks to verify the effectiveness of the proposed LLM-hRIC framework. This scenario was chosen as a foundational proof-of-concept since the resource allocation in multi-cell IAB networks is a well-known difficult problem to handle as it requires the network-wise global information and control. This problem can demonstrate our framework’s strength: using the LLM in the non-RT RIC to provide a globally-aware and coordinated guidance to local RL agents.
@@ III-A Power Allocation in IAB Networks
The IAB networks as illustrated in Fig. 1, consists of a set of macro base stations, denoted as $\mathcal{M}=\{1,\cdots,M\}$, distributed in a network. Each MBS is connected with $N$ small base stations, and $\mathcal{N}_{m}=\{1,\cdots,N\}$ denotes the set of SBSs, connected to the $m$-th MBS. A set of mobile users $\mathcal{K}_{m,n}$, connected with the $n$-th SBS of the $m$-th MBS, follow a Gaussian-Markov mobility model. The SBS allocates fixed power to their connected user. The interference exists among SBSs connected to different MBS that use the same sub-carrier. Both line-of-sight (LoS) and non-line-of-sight (NLoS) path loss characteristics are considered, where the probability of a LoS connection is modeled as $P^{\text{Los}}=\text{exp}(-d/\rho)$, where $d$ represents the distance between MBS and SBS, and $\rho$ is a LoS range constant. Furthermore, the small-scale fading is modeled as independent Nakagami-$m$ fading.
@@ III-B1 LLM-Empowered non-RT RIC
To ensure the reliability of these inputs, a data checking step performs range validation. This process verifies that all metrics are physically plausible, such as ensuring that the ‘number of connected users’ is non-negative and that the ‘average channel gain’ is within a realistic range.
@@ III-C LLM-assisted Three Phases Training for RL
  • Phase 1 (LLM-guided Exploration): The power allocation policy $\boldsymbol{p}_{m}$ is determined by adding noise to the LLM’s initial guidance policy $\boldsymbol{p}_{m}^{o}$, i.e., $\boldsymbol{p}_{m}=\boldsymbol{p}_{m}^{o}+\text{noise}$. This encourages exploration in the vicinity of the LLM’s guidance, allowing the RL agent to collect high-quality experiences and learn a near-optimal initial policy.
  • Phase 2 (Policy Blending): The power allocation policy $\boldsymbol{p}_{m}$ is determined by a weighted combination of the LLM-provided policy $\boldsymbol{p}_{m}^{o}$ and the RL-generated policy $\boldsymbol{p}_{m}^{d}$, i.e., $\boldsymbol{p}_{m}=w\boldsymbol{p}_{m}^{o}+(1-w)\boldsymbol{p}_{m}^{d}$, where the blending coefficient $w$ decays gradually from $1$ to 0 over time. This creates a smooth transition from guided exploration to self-directed optimization.
@@
In each training stage, the power allocation policy interacts with the environment, and the resulting MDP experiences are collected to train the RL agent. This phased approach prevents the instability seen with abrupt policy changes, allowing the agent to efficiently fine-tune its near-optimal policy, which leads to stable convergence and the highest overall throughput.
@@ III-D Experimental Results
Our simulation deploys the Llama-3.1-8B-instruct model [15] in the rApp to generate a power allocation guidance policy on a 2s timescale. For efficient deployment, the model is loaded with bfloat16 precision, with inference parameters of temperature =0.6=0.6 and top_p =0.9=0.9. Meanwhile, a deep deterministic policy gradient (DDPG) agent in the xApp provides a fine-grained power allocation policy for a MBS on a 200ms timescale. The DDPG agent was trained using the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 256.  [^1] Note that the demonstration on a prototype testbed can be done once O1 and A1 interfaces become sufficiently mature in open-source O-RAN platforms. To assess the effectiveness of the proposed framework, we compare it against the following baseline methods: 1) DDPG with linear decay noise for action exploration (DLN), 2) DDPG with cosine decay noise for action exploration (DCN), and 3) equal power allocation (EPA).
Figure 3: LLM and DDPG inference latency with 500 examples.
Fig. 3 shows the LLM and DDPG inference latency with $500$ examples. The LLM inference latency is less than 1.5 seconds, confirming its decision time is well within the operational budget required for the LLM-hRIC ($\leq$ 2s). Moreover, the DDPG inference time is less than $0.07$ ms. This extremely low latency is significantly faster than the 10 ms requirement for the most demanding near-real-time control.
Figure 4: Training curve for proposed with different blending coefficient $w$ and baseline methods over 500 training epochs when $M=3,N=6,W=100\text{Mb},P_{m}^{\text{max}}=44\text{dBm}$, and $\alpha=0.5$.
Fig. 4 shows the training curves of proposed LLM-hRIC framework and baseline methods over $500$ training epochs when the backhaul-access bandwidth partitioning ratio $\alpha=0.5$. The results show that the total throughput increases steadily as training progresses and gradually converges as the models learn effective policies. Notably, the proposed LLM-hRIC framework achieves higher total throughput compared to the baseline methods. This highlights the effectiveness of the LLM-hRIC framework in optimizing resource allocation and improving network performance. Furthermore, the decaying $w$ is critical for performance. A fixed $w=0$ leads to training failure from an abrupt policy shift, while a fixed $w=0.9$ causes performance to stagnate by suppressing the agent’s exploration. In contrast, the proposed decaying $w$ ensures a smooth transition from guided exploration to self-directed optimization, which is necessary for stable convergence to the highest-performing policy.
@@
Fig. 5 illustrates the effect of the backhaul-access bandwidth partitioning ratio on the total throughput. The result shows that the throughput initially increases and then decreases as the allocation ratio $\alpha$ increases. This is because the performance bottleneck exists at the backhaul link for a low $\alpha$, while it is primarily in the access link for large $\alpha$. The proposed LLM-hRIC framework consistently outperforms baseline methods. This is due to the high-quality initial policies, generated by LLM, which enhances exploration efficiency and facilitates better policy discovery by RL. The LLM leverage detailed, and global information to improve coordination among MBSs, resulting in a more efficient resource allocation and higher overall network performance.
@@ IV-B Fine-Tuning of LLMs
Current LLMs are trained for general domains but not optimized for wireless communication tasks, limiting guidance quality in LLM-hRIC. While prompt engineering can partially bridge this gap, its effectiveness is limited by the quality of examples and input token window size. Moreover, the guidance from a general-purpose LLM can sometimes be unstable, making fine-tuning important for adapting to environmental changes and ensuring reliability. Fine-tuning LLMs with offline, domain-specific data is essential to enhance guidance reliability in the non-RT RIC. However, three challenges remain: 1) Collecting or effectively utilizing the limited available data on wireless communication is challenging. 2) Generating high-quality instruction data for fine-tuning is essential. Automated or semi-automated methods for generating diverse, domain-relevant instruction data must be developed to improve the guidance quality of fine-tuned LLMs. 3) A privacy-preserving architecture should be considered such as the combination of federated learning (FL) with differential privacy (DP). The DP adds controlled noise to the shared model updates in FL for a mathematical guarantee that private information on the training data is not leaked.
@@ V Conclusions
This paper proposes the LLM-hRIC framework for O-RAN. In this framework, the LLM-empowered non-RT RIC provides the strategic guidance, while the RL-empowered near-RT RIC explores the guidance with local observation to facilitate near-RT decision-making. The superior performance demonstrates the effectiveness of combining LLM-empowered non-RT RIC and RL-empowered near-RT RIC to tackle the dynamic and complex resource management. Meanwhile, this work highlights the potential of integrating global reasoning and local responsiveness, establishing the LLM-hRIC framework as a foundation for scalable, adaptive, and high-performance solutions in future O-RAN systems. For instance, the framework is extensible to advanced paradigms like the intent-based networking, where the LLM can translate an administrator’s high-level and multi-objective intent into strategic guidance. This guidance is then executed by RL agents with real-time and tactical adaptations, showcasing a powerful synergy for future network automation. Additionally, we identify key challenges for LLM-hRIC, as discussed in Section ‘Future Challenges’, to guide future research and development.